# UI-S1: Advancing GUI Automation via Semi-online Reinforcement Learning

### v1

@@ 1 Introduction
Graphical User Interface (GUI) automation represents a critical frontier in developing AI agents that can interact with digital environments as humans do, driven by advances in multimodal large language models that enable complex reasoning and multi-step task execution (Shen et al., 2023; Hu et al., 2025; Zhang et al., 2025a; Wang et al., 2025a; Tang et al., 2025b; Liu et al., 2025a). This evolution has been accelerated by reinforcement learning techniques that allow agents to improve through trial-and-error learning, guided by task completion signals (Bai et al., 2024; Lu et al., 2025b; Tang et al., 2025a; Ye et al., 2025; Du et al., 2025).
@@ 3 Method
We propose Semi-online RL, a semi-online reinforcement learning framework for training GUI agents that bridges the gap between the stability of offline training and the challenge of online execution. Our approach consists of three key parts. (1) Semi-online rollout (Section 3.2) simulates online interaction dynamics using only offline trajectories; (2) Patch Module (Section 3.3) adaptively recovers the divergence between rollout and expert trajectories; (3) Semi-online Policy Optimization (Appendix 3.4) optimizes agents through a hierarchical reward structure and dual-level advantages.
@@ 3.4 Semi-Online Policy Optimization
Traditional offline RL optimizes only for immediate step-wise accuracy, resulting in multi-turn planning failure. We address this through a hierarchical reward structure and dual-level advantages that capture both immediate and future impacts.
@@
$R_{t}^{i}=\sum_{k=t}^{T}\gamma^{k-t}r_{k}^{i}$ (9)
where $\gamma\in(0,1)$ weights the influence of future consequences on current decisions.
We compute advantages at two levels:
@@
where $R(\tau^{i})$ represents the total trajectory return. We combine these into a unified group-in-group advantage that assigns credit at both global and local scales:
@@
$\mathcal{J}(\theta)=\mathbb{E}_{\begin{subarray}{c}\{\tau^{i}\}_{i=1}^{N}\overset{\mathcal{P}}{\sim}\pi_{\theta_{\text{old}}}(\cdot\mid I)\\ \{o_{i,t}\}_{t=1}^{T}\sim\tau^{i}\end{subarray}}\frac{1}{K}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{k=1}^{|o_{i,t}|}\min\left(\rho(\theta)A(a_{t}^{i}),\text{clip}(\rho(\theta),1\pm\epsilon)A(a_{t}^{i})\right)$ (13)
where the notation $\overset{\mathcal{P}}{\sim}$ indicates trajectories are generated through our Patch Module-enhanced rollout, $K$ is the total number of tokens, $\rho(\theta)=\frac{\pi_{\theta}(o_{i,t,k}|I,o_{i,t,<k})}{\pi_{\theta_{\text{old}}}(o_{i,t,k}|I,o_{i,t,<k})}$ is the importance sampling ratio, and $\beta$ controls the KL penalty strength.
To ensure effective learning with sufficient exploration, we enforce minimum advantage variance: $\sigma(\{A(a_{t}^{i})\})>\eta$, performing dynamic sampling until this diversity threshold is met. In our experiments, we set $\eta=0.3$.
@@ 4.4 Analysis of Training Dynamics
Figure 6 reveals the data scaling performance of Semi-Online RL across different patch configurations. The performance follows an exponential scaling law $y=A+B\cdot e^{C+kx}$, where the scaling coefficient $k$ increases with $\epsilon$ from 1.13-1.13 to 0.73-0.73. This indicates that larger $\epsilon$ values not only improve absolute performance but also enhance data efficiency, enabling more effective learning from each training sample. The improved scaling stems from better utilization of trajectory data, as the Patch Module enables learning from steps that would otherwise be terminated after action mismatches.
@@ 4.5 Ablation Studies
We also conduct ablation studies on training paradigms in Figure 8. Combining SFT with Semi-Online RL outperforms either method alone, achieving 34.0% on AndroidWorld compared to 30.4% for Semi-Online RL only and 21.7% for SFT only. The combined approach also reduces average task completion steps (middle panel), eliminating redundant actions with better planning. Additional ablations (right panel) confirm that both episode-level advantages and maintaining multiple historical images contribute to performance, validating our training setup.
@@ II High-level GUI instruction
A.2 Prompt for Training and Inference
<think> ... </think>
<action> ... </action>
<think>...</think>
<action>{"action": "key", "text": "<value>"}
A.3 Prompt for Thought Generation
<think>
[A coherent reasoning process, reflecting task decomposition,
environmental observation, and iterative decision-making]
</think>
Output Example
<think>
The current task requires checking the order status of
DeepSeek. Access to the official website and locating the login
entry have been completed. Based on the page loading result,
the login form is ready. Authentication information needs
to be filled: the username has already been entered as
"DeepSeek," and now the password must be entered.
</think>
A.4 Patch Module
@@ A.6 Training Details
Our UI-S1-7B is first Supervised Fine-Tuned (SFT) on Qwen2.5VL-7B, trained on data from AndroidControl-Train (Li et al., 2024) and Amex (Chai et al., 2024). The model is then further optimized using Semi-online RL with the thought-free patch mechanism based on the SFT model. The training hyper-parameters are listed in Table 6.
@@ A.7 Training Hours Comparison
A.8 Prompt for GPT-4o to evaluate MiniWob++ Task
A.9 More Cases
@@
A.10 Bad Case
Figure 17: A failed task case in AndroidWorld. The instruction is “Open the file task.html in Downloads in the file manager; when prompted open it with Chrome. Then click the button 5 times, remember the numbers displayed, and enter their product in the form.”. While the model was able to remember the numbers it encountered, it made an error at step $1$, calculating $9\cdot 10\cdot 9\cdot 9\cdot 9$ as $2250$.

### v2

@@ 1 Introduction
Graphical User Interface (GUI) automation represents a critical frontier in developing AI agents that can interact with digital environments as humans do, driven by advances in multimodal large language models that enable complex reasoning and multi-step task execution (Shen et al., 2023; Hu et al., 2025; Zhang et al., 2025a; Wang et al., 2025a; Tang et al., 2025b; Liu et al., 2025a). This evolution has been accelerated by reinforcement learning techniques that allow agents to improve through trial-and-error learning, guided by task completion signals (Bai et al., 2024; Lu et al., 2025b; Tang et al., 2025a; Ye et al., 2025; Du et al., 2025; Zheng et al., 2025).
@@ 3 Method
We propose Semi-online RL, a semi-online reinforcement learning framework for training GUI agents that bridges the gap between the stability of offline training and the challenge of online execution. Our approach consists of three key parts. (1) Semi-online rollout (Section 3.2) simulates online interaction dynamics using only offline trajectories; (2) Patch Module (Section 3.3) adaptively recovers the divergence between rollout and expert trajectories; (3) Semi-online Policy Optimization (Section 3.4) optimizes agents through a hierarchical reward structure and dual-level advantages.
@@ 3.4 Semi-Online Policy Optimization
Traditional offline RL optimizes only for immediate step-wise accuracy, resulting in multi-turn planning failure. We address this through a hierarchical reward structure and dual-level advantages that capture both immediate and future impacts, inspired by GiGPO (Feng et al., 2025).
@@
$R_{t}^{i}=\sum_{k=t}^{t_{\mathrm{end}}}\gamma^{k-t}r_{k}^{i},\quad t_{\mathrm{end}}:=\min\left(\max\left\{k\geq t~\middle|~\forall j\in[t,k],~\text{Matches}(a_{j}^{i},a_{j}^{*})\right\}+1,~T\right)$ (9)
where $\gamma\in(0,1)$ weights the influence of future consequences on current decisions, $t_{\mathrm{end}}$ denotes the final step of the natural (w/o patch) trajectory segment where the predicted actions still match the expert, and $T$ is the index of the last step in the full (w/ patch) trajectory.
@@
where $R(\tau^{i})$ represents the total trajectory return and is computed as $R_{T}^{i}$.
We combine these into a unified advantage that assigns credit at both global and local scales:
@@
$\mathcal{J}(\theta)=\mathbb{E}_{\begin{subarray}{c}\{\tau^{i}\}_{i=1}^{N}\overset{\mathcal{P}}{\sim}\pi_{\theta_{\text{old}}}(\cdot\mid I)\\ \{o_{i,t}\}_{t=1}^{T}\sim\tau^{i}\end{subarray}}\frac{1}{K}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{k=1}^{|o_{i,t}|}\min\left(\rho(\theta)A(a_{t}^{i}),\text{clip}(\rho(\theta),1\pm\epsilon)A(a_{t}^{i})\right)-\beta D_{KL}(\pi_{\theta}||\pi_{\text{ref}})$
where the notation $\overset{\mathcal{P}}{\sim}$ indicates trajectories are generated through our Patch Module-enhanced rollout, $K$ is the total number of tokens, $\rho(\theta)=\frac{\pi_{\theta}(o_{i,t,k}|I,o_{i,t,<k})}{\pi_{\theta_{\text{old}}}(o_{i,t,k}|I,o_{i,t,<k})}$ is the importance sampling ratio, and $\beta$ controls the KL penalty strength. To ensure effective learning with sufficient exploration, we enforce minimum advantage variance: $\sigma(\{A(a_{t}^{i})\})>\eta$, performing dynamic sampling until this diversity threshold is met. In our experiments, we set $\eta=0.3$.
@@ 4.4 Analysis of Training Dynamics
Figure 6 reveals the data scaling performance of Semi-Online RL across different patch configurations. The performance follows an exponential scaling law $y=A+B\cdot e^{C+kx}$, where the scaling coefficient $k$ increases with $\epsilon$ from 1.13-1.13 to 0.73-0.73. This indicates that larger $\epsilon$ values not only improve absolute performance but also enhance data efficiency, enabling more effective learning from each training sample.
@@ 4.5 Ablation Studies
We also conduct ablation studies on training paradigms in Figure 8. Combining SFT with Semi-Online RL outperforms either method alone, achieving 34.0% on AndroidWorld compared to 30.4% for Semi-Online RL only and 21.7% for SFT only. The combined approach also reduces average task completion steps (middle panel), eliminating redundant actions with better planning. Additional ablations (right panel) confirm that both episode-level advantages and maintaining multiple historical images contribute to performance, validating our training setup. More ablations about the hyper-parameter and the reward design are shown in Appendix A.5. We also conduct experiments on 3B and 32B models to investigate the effect of model scale and demonstrate the generalization capability of our method (as shown in Table 6).
@@ II High-level GUI instruction
$t_{\mathrm{end}}$ Last step of a natural trajectory segment
@@
A.2 Patch Module
@@ A.3 SOP
A.4 Model Size Scaling
We conduct experiments on Qwen2.5-VL-3B and Qwen2.5-VL-32B to investigate the impact of model scale in Table 6. Semi-online RL consistently improves performance across all model sizes, demonstrating its strong generalization. The results also show that, as the model size increases, the relative performance improvement diminishes.
Table 6: Performance comparison on different model sizes (3B, 7B, 32B) w/o SFT cold start.
[…]
QwenVL-2.5-3B 3.4 1.4 2.4 5.0 3.7
[…]
QwenVL-2.5-7B 16.8 9.1 13.0 14.9 14.0
[…]
QwenVL-2.5-32B 17.8 10.2 14.0 28.3 21.2
[…]
A.5 Other Ablations
Hyper-parameter
We conduct ablation studies on key hyper-parameters, including $\gamma\in\{0,0.5\}$, $\omega\in\{0,0.5,1\}$, and $\eta\in\{0.1,0.3,0.5\}$, as shown in Table 7. Based on the results, we select $\gamma=0.5$, $\omega=1$, and $\eta=0.3$ for the final training configuration.
Table 7: Ablation on $\gamma$ (future reward discount), $\omega$ (advantage weight), $\eta$ (DAPO threshold) with $\epsilon$ (patch threshold) fixed at 0, data size as 1000 and training epoch as 1. SOP is reported.
$\gamma$ $\omega$ $\eta$ $_{\texttt{PG}} SOP$$_{\texttt{TSR}}$
0.0 0.0 0.1 22.2 11.0
0.0 0.0 0.3 22.3 10.8
0.0 0.0 0.5 21.7 11.4
0.0 0.5 0.1 22.7 12.2
0.0 0.5 0.3 23.3 12.5
0.0 0.5 0.5 22.5 10.2
0.0 1.0 0.1 20.6 11.8
0.0 1.0 0.3 22.2 11.2
0.0 1.0 0.5 22.8 12.1
$\gamma$ $\omega$ $\eta$ $_{\texttt{PG}} SOP$$_{\texttt{TSR}}$
0.5 0.0 0.1 26.8 13.8
0.5 0.0 0.3 26.1 14.6
0.5 0.0 0.5 27.0 13.9
0.5 0.5 0.1 26.9 14.2
0.5 0.5 0.3 27.3 14.0
0.5 0.5 0.5 27.5 14.5
0.5 1.0 0.1 26.5 14.8
0.5 1.0 0.3 27.9 15.4
0.5 1.0 0.5 28.4 14.5
Future reward
We also conduct ablation studies on the choice of $t_{\mathrm{end}}$, defined in Equation 9 (as shown in Table 8). The results show that setting $t_{\mathrm{end}}$ to the last step of a natural trajectory segment achieves better performance compared to using the final step index of the entire trajectory.
Table 8: Ablation on $t_{\mathrm{end}}$ with AndroidWorld success rate reported.
$t_{\mathrm{end}}$ $\epsilon=0$ $\epsilon=1$ $\epsilon=2$ $\epsilon=\infty$
$T$ 25.6 27.9 27.7 27.4
$\begin{aligned} &\min(\max\{k\geq t\mid~\forall j\in[t,k],\\ &~\text{Matches}(a_{j}^{i},a_{j}^{*})\}+1,\,T)\end{aligned}$ 25.6 28.0 28.9 28.4
@@ A.6 Training Details
Our UI-S1-7B is first Supervised Fine-Tuned (SFT) on Qwen2.5VL-7B, trained on data from AndroidControl-Train (Li et al., 2024) and Amex (Chai et al., 2024), then optimized using Semi-online RL with the thought-free patch mechanism. The training parameters are listed in Table 9.
@@
A.7 More Cases
@@
A.8 Bad Case
Figure 17: A failed task case in AndroidWorld. The instruction is “Open the file task.html in Downloads in the file manager; when prompted open it with Chrome. Then click the button 5 times, remember the numbers displayed, and enter their product in the form.”. While the model was able to remember the numbers it encountered, it made an error at step $1$, calculating $9\cdot 10\cdot 9\cdot 5\cdot 5$ as $2250$.
A.9 Prompt for Training and Inference
<think> ... </think>
<action> ... </action>
<think>...</think>
<action>{"action": "key", "text": "<value>"}
A.10 Prompt for Thought Generation
<think>
[A coherent reasoning process, reflecting task decomposition,
environmental observation, and iterative decision-making]
</think>
Output Example
<think>
The current task requires checking the order status of
DeepSeek. Access to the official website and locating the login
entry have been completed. Based on the page loading result,
the login form is ready. Authentication information needs
to be filled: the username has already been entered as
"DeepSeek," and now the password must be entered.
</think>
A.11 Prompt for GPT-4o to evaluate MiniWob++ Task